OpenStudy (kimberly101):

Use equations to solve the following questions In the high jump event Chris leapt 12 centimetres higher than Tim, but their 2 jumps made a total of 3 metres. How high did Chris jump?

OpenStudy (ajprincess):

Let the height of the jump made by Chris = x Then the height of the jump made by Tim will be x-12. x+x-12=3 2x=3+12 2x=15 x=15/2 x=7.5m

OpenStudy (ajprincess):

is that clear?

OpenStudy (kimberly101):

yep thanks

OpenStudy (kimberly101):

the answer says Chris jumped 156 centimetres.

OpenStudy (kimberly101):

????

OpenStudy (campbell_st):

let tim's height = x and chris's height = x + 12 change 3 metres to centimetres x + x+ 12 = 300 2x + 12 = 300 2x = 288 x = 144 chris jumped 144cm , chris jumped 144 + 12 = 156 cm

OpenStudy (campbell_st):

oops tim jumped 144